OpenStudy (anonymous):

Given sin(theta) = 1/2, find cos 2pi - theta. Given sin(theta) = 1/2, find cos 2pi - theta. @Mathematics

OpenStudy (diyadiya):

theta=30degree cos(2pi-30)=cos 30= swrt3/2

OpenStudy (diyadiya):

\[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Oh ok I see. Thank you for breaking it down.